OpenStudy (anonymous):

One of the first mixtures of metals, called amalgams, used by dentists for tooth fillings consisted of 26.0 g of silver, 10.8 g of tin, 2.4 g of copper, and 0.8 g of Zinc. How much silver is in a 25.0 g sample of this amalgam? Please explain your answer clearly and thoroughly so I can understand it! Thanks! :)

OpenStudy (anonymous):

Start by adding up all the metals: 26.0 + 10.8 + 2.4 + 0.8 = 40 You need to find the fraction of silver. So: 26 / 40 = .65 Now that you have the fraction of silver, multiply it times the new amount. .65 * 25g = 16.25g